Find the amount due on a loan of $14,000 at 12%interest compounded annually for 5 years.: Find the amount due on a loan of $14,000 at 12%interest compounded annually for 5 years.
Answer by vleith(1172)

see this --> http://www.moneychimp.com/articles/finworks/fmfutval.htm
FutureValue = Principle(1+r)^n
r = 12% = 0.12
n = 5
Principle = 14000
FV = 14000(1+0.12)^5
FV = 14000(1.12)^5
FV = 14000* 1.7623
FV = 24672.78